### v2

@@ Abstract
In this paper, we propose Conceptual Codebook Learning (CoCoLe), a novel fine-tuning method for vision-language models (VLMs) to address the challenge of improving the generalization capability of VLMs while fine-tuning them on downstream tasks in a few-shot setting. We recognize that visual concepts, such as textures, shapes, and colors are naturally transferable across domains and play a crucial role in generalization tasks. Motivated by this interesting finding, we learn a conceptual codebook consisting of visual concepts as keys and conceptual prompts as values, which serves as a link between the image encoder’s outputs and the text encoder’s inputs. Specifically, for a given image, we leverage the codebook to identify the most relevant conceptual prompts associated with the class embeddings to perform the classification. Additionally, we incorporate a handcrafted concept cache as a regularization to alleviate the overfitting issues in low-shot scenarios. We observe that this conceptual codebook learning method is able to achieve enhanced alignment between visual and linguistic modalities. Extensive experimental results demonstrate that our CoCoLe method remarkably outperforms the existing state-of-the-art methods across various evaluation settings, including base-to-new generalization, cross-dataset evaluation, and domain generalization tasks. Detailed ablation studies further confirm the efficacy of each component in CoCoLe.
@@ 1 Introduction
Pre-trained Vision-Language Models (VLMs), e.g., CLIP [26] and ALIGN [15], have achieved exceptional zero-shot performance in various downstream tasks. Trained on large-scale image-text datasets with contrastive optimization objectives, these models effectively align and embed different modalities into a shared vector space. Despite their impressive performance, adapting these models to diverse downstream tasks remains challenging due to their substantial size. As a result, recent research has concentrated on improving the adaptation of pre-trained VLMs to downstream tasks by adjusting additional parameters while keeping the foundation model fixed. Prompt-tuning methods, e.g. CoOp [42] and ProGrad [43], replace manual prompts with learnable ones to obtain task-specific knowledge, while adapter-based methods utilize extra modules directly on the top of VLMs, such as Clip-adapter [9] and Tip-adapter [38]. These methods have made significant advancements with limited labeled data.
However, we observe that the current fine-tuning methods for CLIP, such as CoOp [42] and CPL [40] demonstrate relatively low performance on fine-grained datasets such as FGVCAircraft [21] (aircraft classification), and UCF101 [31] (action classification). To address the challenge of improving the generalization capability of VLMs in a few-shot settings, in this paper, we propose a novel fine-tuning method called Conceptual Codebook Learning (CoCoLe). Our idea stems from the observation that visual concepts are naturally transferable across domains. As illustrated in Fig. 1, within a single image, there exist multiple distinct visual concepts focusing on different regions. For example, the selected prompts highlight the claws, ears of the koala, and the branches where the koala stands. Moreover, there are similar concepts in images from different classes; for example, the "firetruck", "racer", and "bicycle" classes possess the compound concept of "wheel" in common.
@@
  • We proposed a novel fine-tuning method named CoCoLe for vision-language models to solve the problem of performance degradation on generalization tasks.
@@
  • Extensive experimental results demonstrate the outstanding performance of the proposed CoCoLe compared to existing state-of-the-art methods in base-to-novel generalization, cross-dataset transfer and domain generalization tasks.
Figure 2: Examples and accuracy comparisons on base-to-novel generalization, cross-dataset transfer and domain generalization tasks. S and T represent source and target datasets respectively.
@@ 2.1 Vision-Language Models
In recent times, pre-trained vision-language models (VLMs) has emerged as a notable trend [19, 29, 5, 26]. Capitalizing on tremendous image-text data, these large-scale models can effectively acquire visual representations using contrastive loss, enabling them to grasp both visual and textual semantics and achieve successful modality alignment. Current studies [42, 37] have showcased that by harnessing extensive sets of image-text pairs, VLMs exhibit outstanding performance across a range of downstream visual tasks [13, 6, 16]. For example, Derived through contrastive learning on 400 million online image-text pairs, CLIP [26] demonstrated remarkable zero-shot accuracy on classification tasks. Our method aims to utilize the comprehensive capability of CLIP to perform knowledge-guide fine-tuning for better adaptation to downstream tasks.
@@ 2.2 Prompt Tuning for VLMs
As text input for pre-trained vision-language models, prompt functions as the guidance for the downstream tasks, thereby distilling task-relevant information from inherent knowledge of VLMs [41, 42]. Setting a precedent in this field, CoOp [42] performs end-to-end optimization on the prompt context by a set of learnable vectors but fails to generalize to unseen classes. To address this issue, CoCoOp [41] improved CoOp’s generalization by generating conditional prompts. Further, KgCoOp [36] enhances the generalization by minimizing the discrepancy between learned and handcrafted prompts, and CoPrompt [28] constrains the trainable models by pre-trained ones to avoid overfitting on the downstream task. Meanwhile, there are methods exploring diverse forms of prompts. MaPLe [17] tunes both vision and language branches via a coupling function to induce cross-modal synergy, whereas CPL [40] leverages the powerful generalization of CLIP to build a visual concept cache with a projector to capture multi-level visual features.
@@ 2.3 Visual Concept Learning
Previous research has identified two major approaches to visual concept learning. The first one typically uses manual concept annotations (e.g., colors, textures, and fabric) for the training images [24, 25], while the other method utilizes unsupervised learning to design data-driven concepts [8, 20, 14]. However, these approaches may introduce biases into the learned concepts, thereby limiting their overall performance. Recent studies have sought complementary prompting methods to capture unbiased visual concepts [34, 33]. Notably, CPL [40] first utilizes the capabilities of CLIP [26] to design an unsupervised concept cache. Furthermore, in this work, we introduce a learnable codebook supervised by a handcrafted concept cache, which automatically selects conceptual prompts that are aligned with visual concepts.
@@ 3 Method
In this section, we present the details of our method. First, we provide the background and an overview of our proposed CoCoLe. Second, we delve into the specifics of CoCoLe. Finally, we introduce the training and inference processes of CoCoLe.
@@ 3.1 Background and Overview
CLIP [26] is composed of two encoders: a visual encoder denoted as $E_{v}$ responsible for handling image input $x$, and a text encoder referred to as $E_{t}$ designed to process the corresponding textual prompt $t_{c}$ built as “a photo of $[\text{CLS}]_{c}$", where $[\text{CLS}]_{c}$ represents the word embedding for the class $c$. During training, CLIP learns to optimize the resemblance between the image feature and the prompt embeddings associated with the true label. CoOp [42] further replaces manual prompt construction with learnable prompts by utilizing a collection of $n$ adaptable context vectors $\{[V_{1}],[V_{2}],\cdots,[V_{n}]\}$, each having the same dimension as word embeddings. These vectors are iteratively updated through gradient descent. For a specific class $c$, the respective prompt can be represented as $t_{c}=\{[V_{1}],[V_{2}],\cdots,[V_{n}],[\text{CLS}]_{c}\}$.
@@
In Fig. 3, we present an overview of our proposed CoCoLe method. Figure 3 (a) shows the establishing process of handcrafted concept cache. We first construct a list of text concepts $\Omega_{t}$ that describes major visual concepts. Then we leverage CLIP’s robust text-image correlation capability to discover the image feature $v_{j}$ with Top-$K_{1}$ highest similarity score for each text concept feature $c_{t}^{i}\in C_{t}$. These Top-$K_{1}$matched" features are averaged and stored in the visual concepts cache as keys, with their corresponding text concepts $\omega_{i}\in\Omega_{t}$ as values. Figure 3 (b) shows the handcrafted concept-based prompting process: we first extract the image feature $v$ by $E_{v}$, then use the image feature as the query to find Top-$K_{2}$ similar keys using cosine distance, and finally we utilize the corresponding values with the class embeddings as input for LLM (e.g., GPT [2]) to generate the optimal handcrafted concept-based prompts denoted as $\mathcal{P}^{h}\triangleq\{P_{h_{i}}\}_{i=1}^{N^{C}}$, where $N^{C}$ is the number of class.
Figure 3 (c) presents the training pipeline for CoCoLe. We first extract the visual features $f_{v}$ for a given image $x$ using the visual encoder, Next, we follow (b) to generate the handcrafted concept-based prompts $\mathcal{P}^{h}$ and extract text features by $E_{t}$, denoted as $\mathcal{F}_{h}=E_{t}(\mathcal{P}^{h})$. The visual concepts (keys) and the conceptual prompts (values) in the conceptual codebook are trainable parameters optimized by four loss functions. $\mathcal{L}_{ce}$ is the classification loss adopted to maximize the similarity between image feature $f_{v}$ and the corresponding text features $f_{t}$. $\mathcal{L}_{ma}$ is designed to shorten the distance between the selected keys (Top-$K_{3}$ similar visual concepts) and the image feature $f_{v}$, enabling the keys to learn generalizable concepts. $\mathcal{L}_{cc}$ works as a regularization for diminishing the overfitting problem, ensuring the text features produced by the selected learned prompts do not deviate significantly from those generated by the handcrafted concept-based prompts. Finally, $\mathcal{L}_{or}$ makes the text features of the prompts orthogonal to increase the diversity of the prompts.
@@ 3.2 Conceptual Codebook Learning (CoCoLe)
As such, we propose CoCoLe, as depicted in Fig. 3. The key insight of CoCoLe is a trainable concept codebook, empowering the image to autonomously determine the prompts it should learn based on its inherent concepts. For each training input, only a subset of prompts that align with the current image concepts are chosen and trained individually. The learnable concept codebook stores visual concepts as keys and conceptual prompts as values, which consists of $N$ (key, value) pairs, denoted as $\Psi_{cc}\triangleq\{(\mathbf{V}_{i},\mathbf{P}_{i})\}_{i=1}^{N}$, where $\Psi_{cc}$ denotes the learnable concept codebook, each $\mathbf{V}_{i}\!\in\!\mathbb{R}^{D}$ has the same dimensionality as the image feature $f_{v}$, and each $\mathbf{P}_{i}\!=\![\mathbf{p}_{i}]_{1}\dots[\mathbf{p}_{i}]_{M}\!\in\!\mathbb{R}^{D\times M}$ is composed of $M$ learnable vectors. Denoting the set of learnable visual concepts as $\mathcal{V}\!=\!\{\mathbf{V}_{i}\}_{i=1}^{N}$ and the set of all learnable conceptual prompts as $\mathcal{P}\!=\!\{\mathbf{P}_{i}\}_{i=1}^{N}$. In an optimal scenario, we anticipate that the image itself should determine the prompts to be selected, guided by the concepts it encompasses, in order to steer the prediction process. To this end, given an input image $\mathbf{x}_{j}$, we first obtain its image feature $f_{v_{j}}\!=\!E_{v}(\mathbf{x}_{j})$, where $j$ is the index of the image. Then we calculate the cosine similarity score between $f_{v_{j}}$ and $\mathbf{V}_{i}\in\mathcal{V}$, denoted as,
$S_{c}=\frac{f_{v_{j}}\cdot\mathbf{V}_{i}}{||f_{v_{j}}||~{}||\mathbf{V}_{i}||},$ (1)
Next, we choose the keys with Top-$K_{3}$ cosine similarity score to form set $\mathcal{V}_{j}$, which denotes the subset of Top-$K_{3}$ visual concepts selected from $\mathcal{V}$ specifically for the $j$-th image. We then choose the corresponding conceptual prompts that are paired with these visual concepts, denoted as $\mathcal{P}_{j}\!=\!\{\mathbf{P}_{j_{i}}\}_{i=1}^{K_{3}}$, where $\mathbf{P}_{j_{i}}$ is the $i$-th prompt selected specifically for $\mathbf{x}_{j}$. These prompts are attached to the class name embedding of $\mathbf{x}_{j}$ as illustrated in Fig. 3, and the text input for text encoder can be denoted as, $\mathbf{T}(\mathcal{P}_{j})=\operatorname{concat}(\mathbf{P}_{j_{1}};\dots;\mathbf{P}_{j_{K_{3}}};[\text{CLS}]_{d}),$ where $\operatorname{concat}(\cdot)$ denotes concatenation. Therefore, given test image $\mathbf{x}_{j}$ and the prompts $\mathcal{P}_{j}$ selected according to the concepts of $\mathbf{x}_{j}$, the text feature $f_{t_{j}}$ can be obtained by $f_{t_{j}}\triangleq E_{t}(\mathbf{T}(\mathcal{P}_{j}))$. The probability of predicting the image as class $y_{i}$ is finally computed as:
@@
In Figure 3 (a), following [39], we start by constructing a comprehensive list $\Omega_{t}$ comprising $I=2000$ descriptive text concepts gathered from established visual concept datasets [40, 39]. The descriptions encompass words representing texture, colors, brightness, density etc., and we categorize these descriptions into 50 classes. Illustrations of these terms can be found in Fig. 4. The dictionary is represented as $\Omega_{t}\triangleq\{\omega_{i}\}_{i=1}^{I}$. Adhering to CLIP’s zero-shot setup, we begin by appending $\omega_{i}$ to a manually designed prompt ϕ=italic-$\phi=$The photo is …" to form a concept-specific textual input $\{\pi;\omega_{i}\}$. Consequently, utilizing the text encoder $E_{t}$, we generate text concept features $C_{t}\triangleq\{c_{t}^{i}\}_{i=1}^{I}$, denoted as $c_{t}^{i}=E_{t}({\pi;\omega_{i}})$.
We denote the handcrafted concept cache as $\Phi_{mc}\triangleq\{(key,value)_{i}\}_{i=1}^{I}$. The key and value are the visual concepts and textual concept words respectively. The visual concepts are discovered by leveraging the text concept features $C_{t}$ and the CLIP model, derived from the training images. In the scenario of $H$-shot $D$-class few-shot learning, where there exist $H$ labeled images within each of the $D$ classes. Utilizing the CLIP visual encoder $E_{v}$, we generate their respective image features $V\triangleq\{v_{j}\}_{j=1}^{HD}$, expressed as $v_{j}=E_{v}(x_{j})$. For every text concept feature $c_{t}\in C_{t}$, the similarity score $S_{t}$ is calculated against all visual features in $V$ using the formula $S_{t}=\mathrm{sim}\left(c_{t},v_{j}\right)=c_{t}v_{j}$, where both $c_{t}$ and $v_{j}$ are normalized. Subsequently, we identify Top-$K_{1}$ image features with the highest similarity score and regard their average as the key and its corresponding text concept word $\omega_{t}$ as the associated value, stored within the handcrafted concept cache.
@@
Figure 3 (b) shows the handcrafted concept-based prompting process, we first extract the image feature $f_{v}$ by $E_{v}$, then use the image feature as the query to find Top-$K_{2}$ similar keys using cosine distance, and finally we get the corresponding values (conceptual words). Together with the class name, these concept words are input to an LLM (eg. GPT [2]) to generate optimal handcrafted concept-based prompts. Therefore, our approach addresses the challenge of diminished generalization on downstream tasks by introducing a regularization, ensuring the text features produced by the selected learned prompts are not significantly different from their counterpart generated by the handcrafted concept-based prompts. We enforce this consistency by utilizing the Euclidean distance as a constraint between the text features generated from the hand-crafted concept-based prompts ($f_{t_{d}}^{h}$) and those obtained from selected learned prompts ($f_{t_{d}}^{l}$). While alternative measures such as cosine distance could also serve as constraints, our empirical findings suggest that Euclidean distance yields superior performance. We can represent the consistency constraint as:

### v3

@@ Abstract
In this paper, we propose Conceptual Codebook Learning (CoCoLe), a novel fine-tuning method for vision-language models (VLMs). CoCoLe aims to address the challenge of enhancing the generalization capability of VLMs while adapting them to downstream tasks in a few-shot setting. We recognize that visual concepts like shapes, colors, and textures are inherently transferable across different domains and are essential for generalization tasks. Motivated by this critical finding, we learn a conceptual codebook consisting of visual concepts as keys and conceptual prompts as values, which serves as a link between the image encoder’s outputs and the text encoder’s inputs. Specifically, for a given image, we leverage the codebook to identify the most relevant conceptual prompts associated with the class embeddings to perform the classification. Additionally, we incorporate a handcrafted concept cache as a regularization to alleviate the overfitting issues in low-shot scenarios. This conceptual codebook learning method has been shown to improve the alignment between visual and linguistic modalities. Extensive experimental results demonstrate that our CoCoLe method remarkably outperforms the existing state-of-the-art methods across various evaluation settings, including base-to-new generalization, cross-dataset evaluation, and domain generalization tasks. Detailed ablation studies further confirm the efficacy of each component in CoCoLe.
@@ 1 Introduction
Large-scale pre-trained Vision-Language Models (VLMs), e.g., ALIGN [15] and CLIP [26], have achieved exceptional zero-shot performance in various downstream tasks. These VLMs, trained on massive datasets of image-text pairs with contrastive optimization objectives, effectively align and embed different modalities into a shared vector space. Despite their impressive performance, adapting these models to diverse downstream tasks remains challenging due to their substantial size. As a result, recent research has concentrated on improving the ability of pre-trained VLMs to adapt to downstream tasks by fine-tuning supplementary parameters, while keeping VLMs frozen. Prompt-tuning methods, e.g. CoOp [42] and ProGrad [43], replace manual prompts with learnable ones to obtain task-specific knowledge, while adapter-based methods utilize extra modules directly on the top of VLMs, such as Clip-adapter [9] and Tip-adapter [38]. These methods have made significant advancements with limited labeled data.
However, we find that existing fine-tuning methods for CLIP, including CoOp [42] and CPL [40], exhibit relatively low performance on fine-grained datasets like FGVCAircraft [21] (aircraft classification), and UCF101 [31] (action classification). To address the challenge of enhancing the generalization capability of VLMs in a few-shot settings, in this paper, we propose a novel fine-tuning method called Conceptual Codebook Learning (CoCoLe). Our idea stems from the observation that visual concepts are naturally transferable across domains. As illustrated in Fig. 1, within a single image, there exist multiple distinct visual concepts focusing on different regions. For example, the selected prompts highlight the claws, ears of the koala, and the branches where the koala stands. Moreover, there are similar concepts in images from different classes; for example, the "firetruck", "racer", and "bicycle" classes possess the compound concept of "wheel" in common.
@@
  • We proposed a novel fine-tuning method named CoCoLe for VLMs to solve the problem of performance degradation on generalization tasks.
@@
  • Extensive experimental results demonstrate the outstanding performance of CoCoLe compared to existing state-of-the-art methods in base-to-novel generalization, domain generalization, and cross-dataset transfer tasks.
Figure 2: Illustrations and accuracy comparisons on base-to-novel generalization, cross-dataset transfer and domain generalization tasks. S and T represent source and target datasets respectively.
@@ 2.1 Vision-Language Models
Pre-trained vision-language models (VLMs) have recently emerged as a notable trend [19, 29, 5, 26]. Capitalizing on tremendous image-text data, these large-scale models can effectively acquire visual representations using contrastive loss, enabling them to grasp both visual and textual semantics and achieve successful modality alignment. Current studies [42, 37] have showcased that by harnessing extensive sets of image-text pairs, VLMs exhibit outstanding performance across a range of downstream visual tasks [13, 6, 16]. For example, Derived through contrastive learning on 400 million online image-text pairs, CLIP [26] demonstrated remarkable zero-shot accuracy on classification tasks. Our method aims to utilize the comprehensive capability of CLIP to perform knowledge-guide fine-tuning for better adaptation to downstream tasks.
@@ 2.2 Prompt Tuning for VLMs
As text input for pre-trained vision-language models, prompts function as the guidance for the downstream tasks, extracting task-specific information from the existing knowledge within VLMs [41, 42]. Setting a precedent in this field, CoOp [42] exploits a set of learnable vectors to perform end-to-end optimization on the prompt context but fails to generalize to unseen classes. To address this issue, CoCoOp [41] improved CoOp’s generalization by generating conditional prompts. Further, KgCoOp [36] enhances the generalization by minimizing the discrepancy between learned and handcrafted prompts, and CoPrompt [28] constrains the trainable models by pre-trained ones to avoid the overfitting problem on the downstream task. Meanwhile, there are methods exploring diverse forms of prompts. MaPLe [17] enhances both vision and language components by employing a coupling function to promote cross-modal synergy, whereas CPL [40] leverages the powerful generalization of CLIP to build a visual concept cache with a projector to capture multi-level visual features.
@@ 2.3 Visual Concept Learning
Earlier studies have identified two primary methods for visual concept learning. The first approach generally involves using manual annotations of concepts (e.g., textures, fabrics, and colors) for the training images [24, 25], while the other method utilizes unsupervised learning to design data-driven concepts [8, 20, 14]. However, these approaches may introduce biases into the learned concepts, limiting their overall performance. Recent studies have sought complementary prompting methods to capture unbiased visual concepts [34, 33]. Notably, CPL [40] first utilizes the capabilities of CLIP [26] to design an unsupervised concept cache. Furthermore, in this work, we introduce a learnable codebook supervised by a handcrafted concept cache, which automatically selects conceptual prompts that are aligned with visual concepts.
@@ 3.1 Background and Overview
CLIP [26] consists of two main encoders: the visual encoder $E_{v}$ for processing image inputs $X_{v}$, and the text encoder $E_{t}$ for handling textual prompts $T_{c}$, which are formatted as “a photo of $[\text{CLS}]_{c}$", where $[\text{CLS}]_{c}$ is the word embedding for class $c$. CLIP optimizes the similarity between the image features and the prompt embeddings that correspond to the true labels during training. CoOp [42] takes this a step further by replacing manually constructed prompts with learnable ones. It employs a set of $n$ adaptable context vectors ${[V_{1}],[V_{2}],\cdots,[V_{n}]}$, each with the same dimension as word embeddings. Gradient descent is utilized to optimize the learnable vectors. For a given class $c$, the corresponding prompt is formulated as $T_{c}={[V_{1}],[V_{2}],\cdots,[V_{n}],[\text{CLS}]_{c}}$.
@@
In Fig. 3, we illustrate an overview of our proposed CoCoLe approach. Figure 3 (a) presents the process of constructing the handcrafted concept cache. We begin by compiling a list of text concepts $\Omega_{t}$ that encapsulates key visual concepts. Next, we utilize CLIP’s powerful image-text association capability to identify the image feature $v_{j}$ that has the Top-$K_{1}$ highest similarity scores for each text concept feature $c_{t}^{i}\in C_{t}$. The highest scoring Top-$K_{1}$ features are averaged and saved as keys in the visual concepts cache, paired with their respective text concepts $\omega_{i}\in\Omega_{t}$ as values. In Figure 3 (b), we illustrate the handcrafted concept-based prompting process: initially, we use $E_{v}$ to obtain the image feature $v$. This image feature is then used as a query to identify the Top-$K_{2}$ similar keys based on cosine distance. Subsequently, the associated values, combined with the class embeddings, serve as input for LLM (e.g., GPT [2]) to generate the optimal handcrafted concept-based prompts represented as $\mathcal{P}^{h}\triangleq\{P_{h_{i}}\}_{i=1}^{N^{C}}$, where $N^{C}$ denotes the number of class.
Figure 3 (c) shows the CoCoLe training pipeline. We begin by using the visual encoder to obtain the visual features $f_{v}$ from a given image $x$. We then follow (b) to generate the handcrafted concept-based prompts $\mathcal{P}^{h}$ and obtain text features by $E_{t}$, denoted as $\mathcal{F}_{h}=E_{t}(\mathcal{P}^{h})$. The visual concepts (keys) and the conceptual prompts (values) in the conceptual codebook are trainable parameters optimized by four loss functions. The classification loss $\mathcal{L}_{ce}$ is employed to maximize the alignment between the image feature $f_{v}$ and the related text features $f_{t}$. We utilize the loss function $\mathcal{L}_{ma}$ to minimize the distance between the chosen keys (Top-$K_{3}$ similar visual concepts) and the image feature $f_{v}$, facilitating the learning of generalizable concepts by the keys. $\mathcal{L}_{cc}$ works as a regularization for diminishing the overfitting problem, ensuring the text features produced by the selected learned prompts do not deviate significantly from those generated by the handcrafted concept-based prompts. Finally, $\mathcal{L}_{or}$ ensures that the text features of the prompts are orthogonal to enhance the prompts’ diversity.
@@ 3.2 Conceptual Codebook Learning (CoCoLe)
As such, we propose CoCoLe, as depicted in Fig. 3. The key insight of CoCoLe is a trainable concept codebook, empowering the image to autonomously determine the prompts it should learn based on its inherent concepts. For each training input, only a subset of prompts that align with the current image concepts are chosen and trained individually. The learnable concept codebook stores visual concepts as keys and conceptual prompts as values, comprising $N$ (key, value) pairs, denoted as $\Psi_{cc}\triangleq\{(\mathbf{V}_{i},\mathbf{P}_{i})\}_{i=1}^{N}$, where $\Psi_{cc}$ denotes the learnable concept codebook, each $\mathbf{V}_{i}\!\in\!\mathbb{R}^{D}$ shares the same dimensionality as the image feature $f_{v}$. Additionally, each $\mathbf{P}_{i}\!=\![\mathbf{p}_{i}]_{1}\ldots[\mathbf{p}_{i}]_{M}\!\in\!\mathbb{R}^{D\times M}$ consists of $M$ learnable vectors. We represent the set of learnable visual concepts as $\mathcal{V}\!=\!\{\mathbf{V}_{i}\}_{i=1}^{N}$ and the entire set of learnable conceptual prompts as $\mathcal{P}\!=\!\{\mathbf{P}_{i}\}_{i=1}^{N}$. In an optimal scenario, we anticipate that the image itself should determine the prompts to be selected, guided by the concepts it encompasses, in order to steer the prediction process. To achieve this, for an input image $\mathbf{x}_{j}$, we first extract its image feature $f_{v_{j}}\!=\!E_{v}(\mathbf{x}_{j})$, where $j$ represents image index. Then we calculate the cosine similarity score between $f_{v_{j}}$ and $\mathbf{V}_{i}\in\mathcal{V}$, denoted as, $S_{c}=\frac{f_{v_{j}}\cdot\mathbf{V}_{i}}{||f_{v_{j}}||~{}||\mathbf{V}_{i}||}$.
Next, we choose the keys with Top-$K_{3}$ cosine similarity score to form set $\mathcal{V}_{j}$, representing the subset of Top-$K_{3}$ visual concepts chosen from $\mathcal{V}$ uniquely for the $j$-th image. Then, we select the conceptual prompts that match these visual concepts, represented as $\mathcal{P}_{j}\!=\!\{\mathbf{P}_{j_{i}}\}_{i=1}^{K_{3}}$, where $\mathbf{P}_{j_{i}}$ denotes the $i$-th prompt chosen uniquely for $\mathbf{x}_{j}$. We use these prompts to link to the class name embedding of $\mathbf{x}_{j}$ as shown in Fig. 3, and the input for text encoder can be represented as, $\mathbf{T}(\mathcal{P}_{j})=\operatorname{concat}(\mathbf{P}_{j_{1}};\dots;\mathbf{P}_{j_{K_{3}}};[\text{CLS}]_{d}),$ where $\operatorname{concat}(\cdot)$ signifies concatenation. Thus, for a test image $\mathbf{x}_{j}$ and prompts $\mathcal{P}_{j}$ based on the concepts of $\mathbf{x}_{j}$, the text feature $f_{t_{j}}$ can be obtained by $f_{t_{j}}\triangleq E_{t}(\mathbf{T}(\mathcal{P}_{j}))$. The likelihood of predicting the image as class $y_{i}$ is ultimately determined by:
@@
In Figure 3 (a), inspired by [39], we construct a comprehensive list $\Omega_{t}$ containing $I=2000$ descriptive text concepts sourced from established visual concept datasets [40, 39]. These descriptions encompass terms related to texture, colors, brightness, density, etc., categorized into 50 classes. Examples of these terms are depicted in Fig. 4. The dictionary is defined as $\Omega_{t}\triangleq\{\omega_{i}\}_{i=1}^{I}$. Following CLIP’s zero-shot setup, we start by appending each $\omega_{i}$ to a manually crafted prompt ϕ=italic-$\phi=$The photo is …" to create a concept-specific textual input $\{\pi;\omega_{i}\}$. Subsequently, using the text encoder $E_{t}$, we derive text concept features $C_{t}\triangleq\{c_{t}^{i}\}_{i=1}^{I}$, where each $c_{t}^{i}=E_{t}({\pi;\omega_{i}})$.
We denote the handcrafted concept cache as $\Phi_{mc}\triangleq\{(key,value)_{i}\}_{i=1}^{I}$. The key and value are the visual concepts and textual concept words respectively. The visual concepts are identified by exploiting the text concept features $C_{t}$ extracted from the CLIP model, using training images as a basis. In the context of $H$-shot $D$-class few-shot learning, there are $H$ labeled images available for each of the $D$ classes. Using the CLIP visual encoder $E_{v}$, we obtain their corresponding image features $V\triangleq\{v_{j}\}_{j=1}^{HD}$, where each $v_{j}=E_{v}(x_{j})$. Each text concept feature $c_{t}\in C_{t}$ is matched against all visual features in $V$ using similarity score formula $S_{t}=\mathrm{sim}\left(c_{t},v_{j}\right)=c_{t}v_{j}$, where both $c_{t}$ and $v_{j}$ are normalized. Afterwards, we select the Top-$K_{1}$ image features with the highest similarity scores, computing their average as the key, and associating it with the corresponding text concept word $\omega_{t}$. These pairs are stored within the handcrafted concept cache.
@@
Figure 3 (b) presents the handcrafted concept-based prompting process. Initially, we extract the image feature $f_{v}$ using $E_{v}$. Subsequently, we employ this image feature as the query to retrieve the Top-$K_{2}$ most similar keys based on cosine similarity. Ultimately, we obtain the corresponding values (conceptual words). Together with the class name, these concept words are input to an LLM (eg. GPT [2]) to generate optimal handcrafted concept-based prompts. Therefore, our approach addresses the challenge of diminished generalization on downstream tasks by introducing a regularization, ensuring the text features produced by the selected learned prompts do not differ significantly from their counterpart generated by the handcrafted concept-based prompts. We enforce this consistency by utilizing the Euclidean distance as a constraint between the text features generated from the hand-crafted concept-based prompts ($f_{t_{d}}^{h}$) and those obtained from selected learned prompts ($f_{t_{d}}^{l}$). While alternative measures such as cosine distance could also serve as constraints, our empirical findings suggest that Euclidean distance yields superior performance. We can represent the consistency constraint as: